OpenStudy (anonymous):

Solve: 8x − 2(x + 1) = 7x + 8

OpenStudy (australopithecus):

8x + (-1)2(x + 1) = 7x + 8

OpenStudy (australopithecus):

use distributive property, for what is in the brackets. a(b+c) = a(b) + a(c) Then isolate x on one side of the equation

OpenStudy (australopithecus):

so (-1)2(x+1) = -1(2)(x) + (-1)(2)(1)

OpenStudy (anonymous):

distribute the -2 so you get 8x - 2x - 2 = 7x + 8. Now solve for x. I got -10

OpenStudy (anonymous):

i got -10 too :D

OpenStudy (anonymous):

8x-2x+2=7x+8 8x-2x-7x=8-2 -x=6 x=-6

OpenStudy (anonymous):

how did you get the -10